# Quantum-Informed Machine Learning for Chaotic Systems

### v2

@@ Abstract
Learning the behaviour of chaotic systems remains challenging due to instability in long-term predictions and difficulties in accurately capturing invariant statistical properties. While quantum machine learning offers a promising route to efficiently capture physical properties from high-dimensional data, its practical deployment is hindered by current hardware noise and limited scalability. Here, we introduce a practical quantum-informed machine learning framework for learning partial differential equations, with an application focus on chaotic systems. A quantum circuit Born machine is employed to learn the invariant properties of chaotic dynamical systems, achieving substantial memory efficiency by representing these complex physical statistics with a compact set of trainable circuit parameters. This approach reduces the data storage requirement by over two orders of magnitude compared to the raw simulation data. The resulting statistical quantum-informed prior is then incorporated into a Koopman-based auto-regressive model to address issues such as gradient vanishing or explosion, while maintaining long-term statistical fidelity. The framework is evaluated on three representative systems: the Kuramoto-Sivashinsky equation, two-dimensional Kolmogorov flow and turbulent channel flow. In all cases, the quantum-informed model achieves superior performance compared to its classical counterparts without quantum priors. This hybrid architecture offers a practical route for learning dynamical systems using near-term quantum hardware.
@@ I INTRODUCTION
Modelling chaotic dynamical systems remains one of the most persistent challenges in computational science. Partial differential equations (PDEs) provide the mathematical backbone for describing a wide range of nonlinear, spatiotemporal processes across scientific domains [1, 2, 3] and engineering domains [4]. However, chaotic systems are notoriously sensitive to initial conditions and the floating-point numbers used to compute them [5, 6, 7, 8], making it highly challenging to extract stable, predictive models from data. Modern machine learning techniques often struggle in this regime: while they may fit short-term trajectories, they fail to learn the invariant statistical properties that govern long-term system behaviour. These challenges are compounded in high-dimensional settings, where data are highly nonlinear and contain complex multi-scale spatial-temporal correlations.
@@
Machine learning has seen transformative success in domains such as large language models [9, 10, 11, 12], computer vision [13, 14, 15, 16], and weather forecasting [17, 18, 19, 20], and is increasingly being adopted in scientific disciplines under the umbrella of scientific machine learning [21, 22]. In fluid mechanics in particular, ML has been used to model complex flow phenomena, including wall modeling [23, 24], subgrid-scale turbulence [25, 26], and direct flow field generation [27, 28]. Physics-informed neural networks [29] attempt to inject domain knowledge into the learning process, yet even these models struggle with the long-term stability and generalisation issues that chaotic systems demand. To address this, generative models such as Generative Adversarial Networks (GANs) [30] and operator-learning architectures like DeepONet [31] and Fourier Neural Operators (FNOs) [32] have been proposed. While neural operators offer discretisation invariance and strong representational power for PDE-based systems, they still suffer from error accumulation and prediction divergence over long horizons, particularly in turbulent or chaotic regimes [33, 34]. Recent work, such as Dyslim [35], enhances stability by leveraging invariant statistical measures. However, these methods depend on estimating such measures from trajectory samples, which can be computationally intensive and inaccurate in all forms of chaotic systems. These limitations have prompted exploration into alternative computational paradigms. Quantum machine learning (QML) has emerged as a compelling candidate due to its ability to natively represent and manipulate high-dimensional probability distributions in Hilbert space. Quantum circuits can exploit entanglement and interference to express rich, non-local statistical dependencies using fewer parameters than their classical counterparts, which makes them well-suited for learning invariant measures in chaotic systems, where long-range correlations and multimodal distributions frequently arise [36]. QML and quantum inspired machine learning has already demonstrated potential in fields such as quantum chemistry [37, 38, 39], combinatorial optimisation [40, 41, 42], and generative modelling [43, 44, 45, 46]. However, fully quantum approaches remain constrained by current hardware limitations, such as noise, low qubit counts, and the difficulty of encoding classical physical systems into quantum circuits [47]. These challenges limit the standalone utility of QML in scientific applications today. Instead, hybrid quantum–classical models provide a promising compromise, where quantum submodules are embedded into classical learning pipelines to improve expressivity, data efficiency, and physical fidelity without relying on fault-tolerant quantum computers. Consequently, hybrid architectures have emerged as a promising compromise, enabling practical impact even with noisy intermediate-scale quantum (NISQ) devices. Rather than relying on fully quantum pipelines, these approaches embed quantum modules into classical learning systems to exploit quantum expressivity while retaining classical scalability and training robustness. In quantum chemistry, this hybrid paradigm has proven feasible, notably through quantum mechanical/molecular mechanical (QM/MM) coupling [48, 49, 50], where classical force fields are augmented with quantum corrections. Within such frameworks, techniques like quantum-selected configuration interaction (QSCI) [51] have been employed to enhance accuracy while keeping quantum resource requirements tractable.
Here, we propose a Quantum-Informed Machine Learning (QIML) architecture for learning chaotic dynamical systems. At the core of QIML is a quantum circuit Born machine (QCBM) [52, 53], which is trained on superconducting quantum hardware to learn invariant statistical properties directly from observational data. These properties capture complex, high-dimensional distributions compactly by leveraging entangled quantum states. Unlike standard deep-learning surrogates, such as fully connected or convolutional networks that require large parameter counts to represent such structures, QCBMs operate in exponentially large Hilbert spaces, enabling expressive modelling with as few as 10 to 15 qubits. Crucially, this framework circumvents the data-loading bottleneck common to many quantum machine learning algorithms. The QCBM does not need to encode raw data onto quantum states; instead, it learns a compressed, generative model of the data’s underlying statistical properties. The model is trained with a sample-based Maximum Mean Discrepancy (MMD) loss [54], enabling it to learn the quantum representation without knowing the data distribution. The resulting quantum prior (Q-prior) is a highly memory-efficient “quantum device interpreter” that captures the system’s essential statistics in a compact set of parameters in quantum circuits. The data storage requirement is consistently reduced by over two order of magnitude across the studied systems when using the Q-priors instead of the raw simulation data. The resulting Q-prior is then embedded into a Koopman-based machine learning model, guiding it to maintain long-term statistical fidelity while mitigating autoregressive instabilities such as gradient vanishing or explosion. The QIML framework is implemented as a modular pipeline that integrates quantum-generated priors into a differentiable classical loss [55]. This enables joint optimisation of statistical consistency and dynamical fidelity, including full compatibility with existing superconducting quantum hardware. Even without achieving full quantum advantage [56], the inclusion of Q-priors significantly improves parameter efficiency, prediction stability, and statistical alignment in high-dimensional chaotic regimes. This work demonstrates that QIML provides a practical and scalable hybrid solution for scientific machine learning: one that leverages quantum expressivity while remaining compatible with the practical constraints of current quantum hardware.
In this work, we present the QIML architecture that leverages a quantum-classical hybrid approach for machine learning-based modelling of chaotic dynamical systems. Our QIML scheme is illustrated in Fig. 1. The workflow commences with the data generation stage (Fig. 1a), where we employ a numerical solver on a classical computer to simulate complex chaotic flows, producing the necessary training, validation, and testing datasets. Subsequently, in the quantum machine learning phase (Fig. 1b), a QCBM is implemented on a quantum computer to learn the underlying physical invariant distribution, referred to as the Q-prior. Finally, this Q-prior is used to inform and regularise the classical machine learning model during training (Fig. 1c), guiding it toward long-term statistical consistency with the target dynamical system. The proposed integration demonstrates both practical feasibility and scalability, marking a step towards applying quantum modules in the learning of non-linear partial differential equations. We demonstrate the scalability and robustness of the solution of the QIML framework across a hierarchy of increasingly complex chaotic dynamical systems. First, we evaluate the model on the Kuramoto–Sivashinsky (KS) equation. The solution is discretised into 512 spatial grid points, resulting in a 512-dimensional state vector at each time step. Next, we extend our study to a two-dimensional Kolmogorov flow on a spatial grid, where the QCBM is trained in simulation to extract turbulent invariant measures. Finally, we validate the framework on a cross-section of a three-dimensional turbulent channel flow (TCF). The QCBM is trained on superconducting quantum processors. This progression demonstrates the generality and hardware compatibility of QIML. In all cases, our QIML models outperform classical counterparts lacking Q-prior, particularly in terms of long-term prediction stability, statistical fidelity, and resilience to error accumulation. Crucially, the QCBM acquires these priors directly from raw samples, so the framework remains applicable even when the invariant measure of a high-dimensional chaotic system is unknown. We provide concrete demonstrations of how quantum and classical machine learning can be synergistically combined to address high-dimensional dynamical systems.
II Applications

### v3

# Quantum-informed machine learning for the prediction of chaotic dynamical systems

@@ Abstract
We introduce a quantum-informed machine learning (QIML) framework for the long-term dynamical behavior of high-dimensional chaotic systems. The method combines a one-time, offline-trained quantum generative model with a classical autoregressive predictor for spatiotemporal field generation. The quantum model learns a quantum prior (Q-Prior) that guides the representation of small-scale interactions and improves the modeling of fine-scale dynamics. We evaluate QIML on three representative systems: the Kuramoto–Sivashinsky equation, two-dimensional Kolmogorov flow, and a cross-section of fully developed three-dimensional turbulent channel flow used as a realistic inflow condition. Compared to the classical baseline, QIML yields up to 17.25% improvement in predictive distribution accuracy and a 29.36% improvement in the fidelity of the predicted full energy spectrum. For turbulent channel inflow, the Q-Prior is essential: without it, the model fails to evolve in time, while QIML produces stable, physically consistent forecasts that surpass leading machine learning models for PDEs, including the Fourier Neural Operator and Markov Neural Operator, whose errors diverge. Beyond accuracy, QIML also achieves a memory advantage, compressing multi-megabyte datasets into a kilobyte-scale Q-Prior that captures only the invariant measure needed to guide the classical model, thus circumventing Holevo’s bound by avoiding full data reconstruction. Our findings provide a practical and scalable pathway for integrating the advantages brought by quantum devices into large-scale scientific, engineering modeling and simulation.
@@ I INTRODUCTION
Modeling high-dimensional dynamical systems remains one of the most persistent challenges in computational science. Partial differential equations (PDEs) provide the mathematical backbone for describing a wide range of nonlinear, spatiotemporal processes across scientific [1, 2, 3] and engineering domains [4]. However, high-dimensional systems are notoriously sensitive to initial conditions and the floating-point numbers used to compute them [5, 6, 7, 8], making it highly challenging to extract stable, predictive models from data. Modern machine learning techniques often struggle in this regime: while they may fit short-term trajectories, they fail to learn the invariant statistical properties that govern long-term system behavior. These challenges are compounded in high-dimensional settings, where data are highly nonlinear and contain complex multi-scale spatio-temporal correlations.
@@
Machine learning has seen transformative success in domains such as large language models [9, 10, 11, 12], computer vision [13, 14, 15, 16], and weather forecasting [17, 18, 19, 20], and it is increasingly being adopted in scientific disciplines under the umbrella of scientific machine learning [21, 22]. In fluid mechanics, in particular, ML has been used to model complex flow phenomena, including wall modeling [23, 24], subgrid-scale turbulence [25, 26], and direct flow field generation [27, 28]. Physics-informed neural networks [29] attempt to inject domain knowledge into the learning process, yet even these models struggle with the long-term stability and generalization issues that high-dimensional dynamical systems demand. To address this, generative models such as Generative Adversarial Networks (GAN) [30] and operator-learning architectures like DeepONet [31] and Fourier Neural Operators (FNO) [32] have been proposed. While neural operators offer discretization invariance and strong representational power for PDE-based systems, they still suffer from error accumulation and prediction divergence over long horizons, particularly in turbulent and other chaotic regimes [33, 34]. Recent work, such as Dyslim [35], enhances stability by leveraging invariant statistical measures. However, these methods depend on estimating such measures from trajectory samples, which can be computationally intensive and inaccurate in all forms of chaotic systems, especially in high-dimensional cases. These limitations have prompted exploration into alternative computational paradigms. Quantum machine learning (QML) has emerged as a possible candidate due to its ability to natively represent and manipulate high-dimensional probability distributions in Hilbert space [36]. Quantum circuits can exploit entanglement and interference to express rich, non-local statistical dependencies using fewer parameters than their promising counterparts, which makes them well-suited for capturing invariant measures in high-dimensional dynamical systems, where long-range correlations and multimodal distributions frequently arise [37]. QML and quantum inspired machine learning have already demonstrated potential in fields such as quantum chemistry [38, 39, 40], combinatorial optimization [41, 42, 43], and generative modeling [44, 45, 46, 47]. However, the field is constrained on two fronts: fully quantum approaches are limited by noisy intermediate-scale quantum (NISQ) hardware noise and scalability [48], while quantum-inspired algorithms, being classical simulations, cannot natively leverage crucial quantum effects like entanglement to efficiently represent the complex, non-local correlations found in such systems. These challenges limit the standalone utility of QML in scientific applications today. Instead, hybrid quantum–classical models provide a promising compromise, where quantum submodules work together with classical learning pipelines to improve expressivity, data efficiency, and physical fidelity. In quantum chemistry, this hybrid paradigm has proven feasible, notably through quantum mechanical/molecular mechanical (QM/MM) coupling [49, 50, 51], where classical force fields are augmented with quantum corrections. Within such frameworks, techniques like quantum-selected configuration interaction (QSCI) [52] have been employed to enhance accuracy while keeping quantum resource requirements tractable.
Quantum-enhanced machine learning (QEML) combines the representational richness of quantum models with the scalability of classical learning. By leveraging uniquely quantum properties such as superposition and entanglement, QEML can explore richer feature spaces and capture complex correlations that are challenging for purely classical models. Recent successes in quantum-enhanced drug discovery [53], where hybrid quantum–classical generative models have produced experimentally validated candidates rivaling state-of-the-art classical methods, demonstrate the practical potential of QEML even before full quantum advantage is achieved. Despite these strengths, practical barriers remain. QEML pipelines require repeated quantum–classical communication during training and rely on costly quantum data-embedding and measurement steps, which slow computation and limit accessibility across research institutions. Moreover, most current QEML applications focus on microscopic problems, such as atomic and molecular systems in quantum chemistry, with little exploration of macroscopic phenomena like nonlinear PDEs and turbulent flows, where long-term stability and statistical fidelity pose distinct challenges.
To address these limitations, we propose a Quantum-Informed Machine Learning (QIML) architecture for chaotic systems. The quantum component is a sample-based quantum generator grounded on the quantum circuit Born machine [54, 55], trained once offline on quantum hardware to learn invariant statistical properties directly from observational data. The resulting compact quantum prior (Q-Prior) can be deployed broadly and efficiently without ongoing quantum–classical iteration, eliminating the need for repeated access to quantum hardware during training. This design is specifically tailored to learning the dynamics of nonlinear PDEs. These properties capture complex, high-dimensional distributions compactly by leveraging entangled quantum states. Unlike standard deep-learning surrogates, such as fully connected or convolutional networks that require large parameter counts to represent such structures, the quantum generators operate in exponentially large Hilbert spaces, enabling expressive modeling with as few as 10 to 15 qubits for the systems in this paper. Crucially, this framework circumvents the data-loading bottleneck common to many QML algorithms. The generator does not need to encode raw data onto quantum states; instead, it learns a compressed, generative model of the data’s underlying statistical properties. The model is trained with a sample-based Maximum Mean Discrepancy (MMD) loss [56], enabling it to learn the quantum representation without knowing the data distribution. After the training of the quantum generator, the resulting Q-prior is then integrated into an autoregressive, unitary-enhanced Koopman-based machine learning framework [57]. This architecture is powerful for learning dynamical systems as it can linearize nonlinear dynamics in a lifted observable space, making it effective for capturing coherent structures [58, 59]. However, for high-dimensional chaotic systems, its ability to maintain predictive stability remains a significant challenge, often failing to preserve the systems’ physical invariant measures over extended rollouts (i.e. recursive 1-in-1-out predictions where each forecasted state is fed back as input for the next step). This is particularly true when modeling data for which no closed-form governing equation is known—as is the case for a two-dimensional cross-section extracted from a three-dimensional turbulent domain. We address this limitation by embedding the Q-prior as a differentiable constraint directly within the loss function. This composite loss, which penalizes divergence from the Q-prior, imposes a powerful statistical regularization on the evolution of the Koopman operator, guiding it to learn a physically consistent and stable long-term dynamic. We evaluate our QIML framework against its classical counterpart without Q-informed guidance, as well as two leading machine learning models: the FNO for general PDE-based systems and the Markov Neural Operator (MNO) specifically designed for chaotic dynamics. Our QIML framework outperforms all baseline models, which fail to sustain stable rollouts over time, leading to divergence and loss of physical consistency.
The remainder of this paper is structured as follows. In Sec. II, we detail the architecture of our QIML framework, including the Koopman operator formulation and the integration of the quantum prior. In Sec. III, we present the evaluation of our framework on three canonical chaotic systems. A direct comparative analysis of QIML against state-of-the-art classical machine learning baselines is provided in Sec. IV. We then discuss the profound memory efficiency of the Q-Prior in Sec. V. Finally, we conclude in Sec. VIII and offer a discussion on the limitations and future outlook of this work in Sec. VII.
II Quantum-informed machine learning framework
In this section, we present the QIML architecture that leverages a quantum-classical hybrid approach for machine learning-based modeling of high-dimensional dynamical systems. Our QIML scheme is illustrated in Fig. 1. The workflow commences with the data generation stage (Fig. 1a), where we employ a numerical solver on a classical computer to simulate complex high-dimensional flows, producing the necessary training, validation, and testing datasets. Subsequently, in the quantum machine learning phase (Fig. 1b), a quantum generator is implemented on a quantum computer to learn the underlying physical invariant distribution, referred to as the Q-Prior. Finally, this Q-Prior is used to inform and regularize the classical machine learning model during training (Fig. 1c), guiding it toward long-term statistical consistency with the target dynamical system. We begin by introducing the machine learning component, which is based on a Koopman operator formulation. The Koopman operator offers a linear representation of nonlinear dynamical systems by acting on observables instead of directly on the state variables. In this framework, the evolution of the system is captured not in the original state space, but in a lifted function space where the dynamics become linear. Consider a discrete-time dynamical system:
where with denotes the dimension of the state space at time , and is a generally nonlinear transition function. To enable auto-regressive prediction and long-term rollouts, we aim to find a mapping from the high-dimensional space to a latent space where a linear operator governs the time evolution. The measure-preserving property in the latent space is particularly important for capturing the statistical properties and invariant measure of high-dimensional dynamical systems. The Koopman operator acts on observables such that:
While is linear, our implementations seek finite-dimensional approximations by learning a set of observables whose evolution under is closed and linear. Next, we approximate the Koopman operator using a neural network; we define an encoder , with denotes the latent dimension, that maps the system state into a latent (feature) space, and a decoder that reconstructs the original state. Unlike generic autoencoders primarily focused on data reconstruction, our Koopman operator neural network is specifically designed to learn a linear, norm-preserving evolution in the latent space, which is crucial for maintaining the long-term stability and physical fidelity of high-dimensional dynamical systems. Our detailed autoencoder-decoder architecture description can be found in Appendix G for 1D and 2D cases. The dynamics in the latent space are modeled by a linear operator , such that:
where is the encoded latent state and is the reconstructed state. To preserve the long-term statistical behavior of the system and approximate the spectral properties of a measure-preserving dynamical system, we constrain the operator to be unitary, i.e., [60, 57, 61]. This ensures that the Koopman dynamics in the latent space are norm-preserving and stable under long-term rollouts. In this work, unitarity is imposed by adding a regularization term to the loss function of the form . The Koopman operator evolves the system in a latent Hilbert space of encoded observables, enabling a linear and spectrally coherent representation of the dynamics. However, the preservation of long-term statistical properties in this latent space does not automatically guarantee consistency with the statistical structure of the original physical space.
To address this, we integrate a sample-based quantum generator, which approximates the invariant distribution Q-Prior of the system. This quantum-learned prior is used as a statistical regularizer and physically informed guidance during the classical machine learning training, guiding the model to produce predictions that are not only temporally coherent but also statistically consistent with the system’s long-term behavior (See Method C). This integration is achieved by incorporating the Q-Prior into a composite loss function that combines reconstruction fidelity with statistical alignment metrics, which is shown below.
While the Koopman operator is introduced using a generic state variable, in this work, we focus on predicting physical states . We use the Koopman operator to predict the next time frame of the physical states . The training of our model is guided by a composite loss function. We define its components below.
First, to ensure predictive accuracy, we use a standard reconstruction loss which measures the mean squared error between the predicted state and the ground truth :
Second, we include the unitary regularization term as previously discussed:
where is the identity matrix and denotes the Frobenius norm. Finally, we incorporate the Q-prior, , as a statistical constraint. This is achieved through two complementary metrics. The Kullback-Leibler (KL) divergence captures low-order, information-theoretic discrepancies between the empirical distribution of predicted states , and the Q-Prior:
To align higher-order statistical moments, we also include a MMD term :
These individual components are combined into a single training objective, where the Q-Prior terms are weighted by empirically selected hyperparameters and :
The full QIML framework is demonstrated in Algorithm 1. All models are implemented in PyTorch with fully differentiable integration of Q-Priors. Architectural details for predicting 1D and 2D systems are provided in Appendix G.
II.1 Invariant measure for high-dimensional dynamical systems
The concept of an invariant measure originates from the field of ergodic theory and has become a fundamental tool in the study of dynamical systems. First formalized in the context of measure-preserving transformations by Birkhoff and von Neumann in the early 20th century, invariant measures capture the long-term statistical behavior of systems evolving under deterministic rules [62, 63].
which means that the total probability of the system being in a region remains the same after one step of evolution. This property reflects a kind of statistical balance: although individual trajectories may change, the overall distribution stays stable. In our work, we aim to capture this long-term behavior by ensuring that the model’s predictions remain consistent with a reference distribution, which we learn using a quantum generator.
In turbulent flows, such invariant measures typically manifest as conserved distributions, such as velocity probability, turbulent kinetic energy, or energy spectra that encode physical long-term steady states. Capturing these invariant features is essential for ensuring robustness in extrapolative prediction regimes, where a classical machine learning method often suffers from gradient explosion, mode collapse, or long-term error accumulation [64].
II.2 The quantum generator to learn the Q-Prior
We develop a quantum-informed learning component in which a quantum generator is used to directly learn invariant statistical properties from high-dimensional physical observations. The generator defines a parameterized quantum state on an -qubit register, which serves as a generative model over a discretised spatial domain. It does not encode a specific state (e.g., a velocity field) into the circuit. Instead, by observing many snapshots of the system, it learns to reproduce its underlying invariant probability distribution. Specifically, each computational basis state here is mapped bijectively to a spatial coordinate in the simulation grid. The corresponding amplitude is optimized so that its squared magnitude approximates the normalized velocity magnitude at that point. Although the circuit does not encode classical inputs directly via amplitude encoding, it constructs a quantum state whose Born distribution approximates the invariant measure Q-Prior of the system. The resulting Born distribution is given by:
where is a parameterized unitary circuit composed of L layers, each consisting of single-qubit rotations and entangling operations. Each computational basis state represents a discrete spatial point, and the quantum amplitudes are trained to reflect the empirical distribution of local velocity magnitudes. This provides a compact, data-driven mapping from physical observables into the Hilbert space. The exponential size of this space, combined with the use of entangling gates, provides a natural mechanism to encode the complex, non-local correlations inherent in chaotic systems. This feature provides a quantum advantage rooted in the ability of quantum systems to represent high-dimensional classical data with exponential efficiency [65]. This enables it to learn the required statistical properties using exponentially fewer parameters than would be required by a classical model [66]. Furthermore, by leveraging the vast Hilbert space and entanglement, the quantum circuit can capture the complex, non-local correlations inherent in such systems efficiently [67]. Further details can be found in section VI and Appendix A.
As shown in Algorithm 2, the quantum circuit is trained as a quantum generative model. Each computational basis state is mapped uniquely to a spatial location in the discretised velocity field, and its target value is the corresponding velocity magnitude . Because the underlying probability distribution over spatial locations is unknown, we adopt a sample–based training strategy that minimises the MMD loss:
Here, denotes the batch of empirical samples drawn from the reference velocity field, while are samples generated by the quantum circuits. Because the MMD objective is expressed solely through kernel evaluations on finite sample sets, optimization of the generator requires no closed-form expression for the target probability density. This sample-based objective is particularly advantageous for high-dimensional chaotic flows, where the invariant measure is available only as a time-ordered set of simulation snapshots rather than an explicit probability density. By aligning the quantum generator’s samples with these numerical trajectories, the model infers the underlying measure without ever needing a closed-form expression.
The trained quantum distribution thus can serve as a data-driven prior capturing the invariant measure of the underlying physical process. Unlike classical regularizers based on fixed heuristics, the quantum generator constructs this Q-Prior through optimization on empirical data.
All quantum circuits are implemented using hardware-efficient structures, which are detailed in Appendix B. Training is carried out on noiseless simulators and on superconducting quantum hardware, with circuit transpilation, error mitigation, and device-specific calibration performed using the Qiskit runtime. Further details and the optimization methods are provided in Appendices A and B.
II.3 Quantum hardware implementation
The sample-based quantum generator model was trained using both emulators on classical computers and real superconducting quantum processors. A classical emulator here is a program running on a conventional computer that simulates the mathematical operations of an ideal, noise-free quantum device. In this work, these emulations were supported by the PennyLane and Qiskit software packages. The practical constraints of current NISQ-era hardware, which require cryogenic operating temperatures and are subject to frequent maintenance and recalibration, make continuous, large-scale training across numerous datasets challenging. For this reason, we adopted a strategic approach here. The first two systems (Kuramoto-Sivashinsky and Kolmogorov flow) were studied on the classical emulator to establish the Q-Prior’s efficacy in an ideal, noise-free setting. We then reserved the use of the real quantum hardware for the most complex and challenging dataset, which is the turbulent channel flow, precisely because this is the case where leading classical methods fail, providing the most stringent test for the QIML framework’s practical utility and robustness under realistic noise conditions. To ensure effective execution on near-term quantum hardware and to manage the impact of noise on the quality of Q-Priors, this study constrains the total number of trainable parameters and adopts conservative circuit configurations. Specifically, the number of qubits is 10–15 on the emulator and 10 on the hardware, with circuit depths ranging from 3 to 8 layers, resulting in fewer than 300 trainable parameters in total (see Appendix B table 4). Hardware experiments were mainly conducted on IQM’s 20-qubit superconducting chip (Garnet) [68, 69], using qubits 1–10 as the active subset, supported by Qiskit and IQM-Qiskit. Further discussion on hardware implementation and sensitivity is provided in Appendix B. To ensure scalability, the quantum emulation and hardware backend were integrated into a GPU-accelerated classical training loop on NVIDIA A100 nodes provided by the BEAST GPU cluster (Leibniz Supercomputing Centre). Gradient evaluations, loss computations, and optimizer updates were performed classically and synchronised with the quantum hardware at each iteration. This setup constitutes an applied demonstration of embedding quantum devices into scientific machine learning pipelines.
The employed quantum circuit architecture consisted of alternating layers of parameterized rotations and hardware-native entangling gates. Importantly, the same circuit structure was used across all experiments without dataset-specific tailoring, ensuring general applicability. Although performance could potentially be improved through customised ansatz design, our results demonstrate that the proposed quantum-informed framework is compatible with any hardware-efficient quantum circuit. This adaptability enables the integration of diverse circuit architectures without modifying the overall training pipeline, suggesting broad utility across problem domains and quantum backends.
The Q-prior is generated via a one-time, offline training process. This training consists of a total of 50 epochs. Following the completion of this training, the quantum circuit is not utilized during the main classical model’s training loop. For each of the 50 training epochs, the circuit was sampled with 20,000 shots to ensure a high-fidelity statistical distribution. Readout errors were mitigated using IQM’s native transpiler and the M3 measurement mitigation toolkit [70]. While we used this number of shots to ensure maximum stability and demonstrate the framework’s potential, this is not a strict requirement for its effectiveness. In separate tests, we reduced the measurement count to 8,000 shots and found the impact on the final model’s performance was less than 5%, indicating that the Q-Prior remains effective. Nevertheless, to present the best possible outcomes in this work, all results and figures are based on data obtained from 20,000 shots per iteration. optimization was performed using L-BFGS [71], Adam [72], and COBYLA [73], selected based on convergence stability and noise robustness. To prevent mode collapse, such as excessive sampling concentration on the state, runtime monitoring and redundancy checks were incorporated into the training loop. Despite operating with only 10 to 15 qubits, the Q-Prior successfully captured the structure of invariant distributions over a - to -dimensional Hilbert space, demonstrating the effectiveness of quantum priors in modeling high-dimensional systems. Further discussions of the error mitigations, experiment results, and ablations are provided in Appendices B and G.
III Results
III.1 Numerical set up